# On the degrees of freedom count on singular phase space submanifolds

### v3

@@ 5 Discussion
Formally, one can make the $m=0$ equations similar to the massive ones by imposing the Lorenz gauge of $\partial_{\mu}A^{\mu}=0$. But it’s not possible to do so in a model with an effective mass parametrised by a scalar field (12). It would require a condition of $\partial_{\mu}(\psi A^{\mu})=0$ which then pretty much depends on how the scalar field behaves in the regime of its small values. Even more intricate catastrophes can happen if the mass depends on a more elaborate object than a mere scalar field.
As an old and pathological but rather curious example, one can recall the vector inflation [19]. Normally, vector fields decay in an expanding universe. So, we gave them a tachyonic mass. Of course, it immediately implies the problem of longitudinal ghost [20], even though one can have doubts about how problematic it is [21]. Still, even if the inflationary behaviour can be reliably described, there must also be a moment in time when the effective mass goes through zero, for the tachyonic inflaton should probably become a normal matter component later on. This is a domain in which the model does not possess a well-defined number of degrees of freedom, and there is no surprise that it does not go smoothly [20, 22].
At the same time, there is also an outright problem in this approach which, strangely enough, was noticed only later [22, 23]. Namely, the effective mass got its main contribution from a non-minimal coupling of $RA^{\mu}A_{\mu}$ type. Unless we go for Palatini formulation, the scalar curvature contains second derivatives of the metric. It means that the gravitational equations feature the second time derivatives of all the vector field components including $A_{0}$. In other words, all four vector field components appear to be dynamical [^5] Actually, there are many vector fields in vector inflation. They cannot all obtain dynamical temporal components via the gravitational dynamics only. However, one combination of those does become dynamical as can be seen via a conformal transformation to Einstein frame [22]., and that can hardly be any stable. Moreover, the very inflationary background itself presents a strong coupling for this extra mode, for it has $A_{0}=0$ and the unwanted acceleration comes in the combination of $A_{0}\ddot{A}_{0}$.
@@ 5.1 On the $R+R^{2}$ and pure $R^{2}$ gravities again
As we see, presence of gravity makes the pathologies even more complicated, even when it concerns the usual metric approach to such theories. Coming back to that, I confirm that the $f(R)=R^{2}$ gravity linearised around Minkowski spacetime has zero dynamical degrees of freedom [1], at least when taken at the face value. It happens due to an accidental gauge symmetry which obviously emerges at the point of $f^{\prime}=0$. The vanishing of tensorial polarisations is then a straightforward consequence of the fact that these modes have fully become pure gauge. What used to be a usual gravitational wave is no longer in the action at all.
@@
To sum it up, the ”pure $R^{2}$ gravity”, i.e. $f(R)=R^{2}$, and the ”full $R^{2}$ gravity”, i.e. $f(R)=R+R^{2}$, are genuinely different from each other if to think of their behaviour around Minkowki space. This space is a singular locus for the former and is completely regular for the latter. Otherwise, they are very similar. Generically, both have three dynamical degrees of freedom and four more physical modes which are constrained, and both have singular loci in their phase spaces, with the only difference in that it is about the points with $R=0$ for the former, and about some non-zero value of $R$ for the latter.
@@
Talks of strong coupling scales [1] are irrelevant. These effects of strong coupling which we face at $f^{\prime}=0$ are not just about an interaction being too strong for being taken into account as a small perturbation. It is about mathematical ill-posedness of an initial value problem for classical equations. Roughly speaking, it is an infinitely strong coupling. In particular, one can introduce a small non-zero cosmological constant to the pure $R^{2}$ gravity and study perturbations around the corresponding de Sitter solution. Those will have three dynamical modes and require the corresponding Cauchy data. Moreover, locally the full set of exact solutions, as long as they never touch the value of $R=0$ in a domain under consideration, features the same freedom of varying the initial data.
@@ 5.2 On modified teleparallel degrees of freedom
The situation in $f(\mathbb{T})$ is very unclear. The available Hamiltonian analyses [24, 25, 26] do not agree with each other, not even in the number of degrees of freedom. In 4D, and on top of the usual two polarisations, the old claim [24] is that there are three new degrees of freedom (that is five in total). The analysis is not very clean, and even a much more accurate recent account [26] confirming this result does not offer any reliable analysis of when the jumps in degrees of freedom occur. That would require a detailed exposition of what happens with the numbers of constraints and the ranks of their Poisson brackets algebra throughout the phase space. Another Hamiltonian analysis [25], which claimed only one extra dynamical mode, had for sure committed at least one essential mistake [26].
Given this fact, it is really amazing to look at what we have seen in perturbations. Around the trivial Minkowski background of a unit matrix tetrad, there are no new modes at all, due to an obvious reason: an accidental restoration of the full local Lorentz symmetry takes place. Much more surprising is that no new dynamical modes were found around simple cosmological backgrounds either [27], not even when spatial curvature was switched on [28]. The local Lorentz symmetry is no longer there, but the new physical modes are constrained. At the same time, an investigation of a non-trivial tetrad solution corresponding to Minkowski metric had found one extra dynamical mode [9], though with a somewhat restricted Cauchy data freedom similar to an example of the variable mass Proca field from above. In other words, it rather confirms the results of Ref. [25]
One possible explanation for this strange coincidence is almost obvious. The thing is that the non-trivial tetrad [9], even though being quite non-symmetric, did preserve the condition of ${\mathbb{T}}=0$, while the main mistake of the Hamiltonian analysis [25] was in missing the contributions to Poisson brackets coming from spatial derivatives of the auxiliary scalar field corresponding to $\mathbb{T}$. However, this explanation is then very intriguing in itself. In almost every cosmological setting, away from bounces and the like, the torsion scalar has got a time-like gradient. Then $\mathbb{T}$ can be used precisely as a time variable, and the analysis of Ref. [25] seems to apply perfectly well. If true, it might mean that the model enjoys a preferred foliation, similar to the story of cuscuton fields [29].
@@ 5.3 Conclusions
When exploring a new theoretical idea, we usually start from a Lagrangian density and the variational principle. It gives us very convenient tools for investigation. However, the Hamiltonian formalism is also very important for a variety of reasons. It has a very nice symplectic geometry in it which is good for quantising if we understand anything about quantum physics. And even safely staying inside the classical framework, we happily get a first-order system of time evolution equations. It’s crucial for the whole range of theoretical research, from numerical implementations if needed to rather abstract analyses of mathematical features.
In the regular situations, the Hamiltonian formalism is very well understood, but unfortunately it is very often used as just a recipe from a textbook. However, especially when going for teleparallel geometries, we often encounter ill-posed systems of equations. They might very well be a dead end, but we anyway need to understand what happens with them and how. In my opinion, it is a very interesting and important task, to learn more about it and to see what the Hamiltonian analysis might tell us about the properties of such Lagrangian systems.

### v5

@@ 5 Discussion
Formally, one can make the massless equations similar to the massive ones by imposing the Lorenz gauge of $\partial_{\mu}A^{\mu}=0$. But it’s not possible to do so in a model with an effective mass parametrised by a scalar field (12). It would require a condition of $\partial_{\mu}(\psi A^{\mu})=0$ which then pretty much depends on how the scalar field behaves in the regime of its small values. Even more intricate catastrophes can happen if the mass depends on a more elaborate object than a mere scalar field.
As an old and pathological but rather curious example, one can recall the vector inflation [26]. Normally, vector fields decay in an expanding universe. So, we gave them a tachyonic mass. Of course, it immediately implies the problem of longitudinal ghosts [27], even though one can have doubts about how problematic they really are [28]. Still, even if the inflationary behaviour can be reliably described, there must also be a moment in time when the effective mass goes through zero, for the tachyonic inflaton should probably become a normal matter component later on. This is a domain in which the model does not possess a well-defined number of degrees of freedom, and there is no surprise that it does not go smoothly [27, 29].
At the same time, there is also an outright problem in this approach which, strangely enough, was noticed only later [29, 30]. Namely, the effective mass got its main contribution from a non-minimal coupling of $RA^{\mu}A_{\mu}$ type. Unless we go for Palatini formulation, the scalar curvature contains second derivatives of the metric. It means that the gravitational equations feature the second time derivatives of all the vector field components including $A_{0}$. In other words, all four vector field components appear to be dynamical [^7] Actually, there are many vector fields in vector inflation. They cannot all obtain dynamical temporal components via the gravitational dynamics only. However, one combination of those does become dynamical as can be seen via a conformal transformation to the Einstein frame [29]., and that can hardly be any stable. Moreover, the very inflationary background itself presents a strong coupling for this extra mode, for it has $A_{0}=0$ and the unwanted acceleration comes in the combination of $A_{0}\ddot{A}_{0}$.
To summarise the landscape of the strong coupling issues, they generically tend to appear when the constraints have a complicated non-linear structure [24, 25]. The most obvious problematic loci are extremal points of non-linear functions of kinetic terms in the action functional, like the locus of $f^{\prime}=0$ in $f(R)$ gravity. Another common example is $f^{\prime\prime}=0$ if the argument of the function contains an essential quantity in the shape of a total (Levi-Civita) divergence. It applies both to $f(R)$ and to non-linear extensions of teleparallel gravities. And the whole field of modified teleparallel gravity is even more problematic in general, due to gauge symmetries of GR-equivalent options being broken in a very feeble manner, with rich but severely unstable sets of remnant symmetries [2]. Then, even without gravity, if a locus of zero mass restores a broken gauge invariance, it is a singular option. In gravity theories, it applies to massive gravity of variable mass and to symmetry-breaking non-minimal couplings of vector fields. Finally, one can intentionally produce an ill-defined Cauchy problem [4, 5] by changes of variables of a singular type, like $x\leftrightarrow x^{3}$, inside the action functional.
@@ 5.1 On the $R+R^{2}$ and pure $R^{2}$ gravities again
As we see, presence of gravity makes many pathologies more complicated, even when it concerns the usual metric approach to such theories. The recent paper [1] considered a very simple case of pure $R^{2}$ gravity around its singular locus of Minkowski spacetime. Strangely enough, they present two different methods of calculating the number of degrees of freedom which yield contradictory results. They finally stick to the correct answer of no dynamical modes at all, however do not provide a proper explanation for what has happened in the alternative approach. Moreover, they then continue with an absolutely wrong statement that the absence of dynamical degrees of freedom extends beyond the linear order and to the full $R+R^{2}$ gravity, too.
Coming back to that, I confirm that the $f(R)=R^{2}$ gravity linearised around Minkowski spacetime has zero dynamical degrees of freedom [1], at least when taken at the face value. It happens due to an accidental gauge symmetry which obviously emerges at the point of $f^{\prime}=0$. The vanishing of tensorial polarisations is then a straightforward consequence of the fact that these modes have fully become pure gauge. What used to be a usual gravitational wave is no longer in the action at all.
@@
The dynamical mode(s) seen in the ”covariant approach” [1] are indeed due to the ambiguity of the spin projector formalism (30), or what they called an incomplete gauge fixing. However, it can only explain one of the two dynamical modes of the 4th order equation they had got. Another mode comes from the mistake of fixing a gauge inside the action functional. Unfortunately, these problems of gauge-fixed Lagrangians are rarely discussed, though there are nice exceptions, for example in the realm of cosmological perturbations [31].
To sum it up, the ”pure $R^{2}$ gravity”, i.e. $f(R)=R^{2}$, and the ”full $R^{2}$ gravity”, i.e. $f(R)=R+R^{2}$, are genuinely different from each other if to think of their behaviour around Minkowki space. This space is a singular locus for the former and is completely regular for the latter. Otherwise, they are very similar. Generically, they both have three dynamical degrees of freedom and four more physical modes which are constrained, and both have singular loci in their phase spaces, with the only difference in that it is about the points with $R=0$ for the former, and about some non-zero value of $R$ for the latter.
@@
Talks of strong coupling scales [1] are irrelevant. These effects of strong coupling which we face at $f^{\prime}=0$ are not just about an interaction being too strong for being taken into account as a small perturbation. It is about mathematical ill-posedness of an initial value problem for classical equations. Roughly speaking, it is an infinitely strong coupling. In particular, one can introduce a small non-zero cosmological constant to the pure $R^{2}$ gravity and study perturbations around the corresponding de Sitter solution. Those will have three dynamical modes and require the corresponding Cauchy data. Moreover, locally the full set of exact solutions, as long as they never touch the value of $R=0$ in a domain under consideration, features the same freedom of varying the initial data. The disappearance of each and every dynamical mode is an artifact of using the sick background. Of course, it can also be seen in a purely quantum-field-theory language [32].
@@ 5.2 On modified teleparallel degrees of freedom
The situation in $f(\mathbb{T})$ is very unclear. The available Hamiltonian analyses [33, 34, 35] do not agree with each other, not even in the number of degrees of freedom. In 4D, and on top of the usual two polarisations, the old claim [33] is that there are three new degrees of freedom (that is five in total). The analysis is not very clean, and even a much more accurate recent account [35] confirming this result does not offer any reliable analysis of when the jumps in degrees of freedom occur. That would require a detailed exposition of what happens with the numbers of constraints and the ranks of their Poisson brackets algebra throughout the phase space which is by far not an easy task. Another Hamiltonian analysis [34], which claimed only one extra dynamical mode, had for sure committed at least one essential mistake [35].
Given this fact, it is really amazing to look at what we have seen in perturbations. Around the trivial Minkowski background of a unit matrix tetrad, there are no new modes at all, due to an obvious reason: an accidental restoration of the full local Lorentz symmetry takes place. Much more surprising is that no new dynamical modes were found around simple cosmological backgrounds either [36], not even when spatial curvature was switched on [37]. The local Lorentz symmetry is no longer there, but the new physical modes are constrained. At the same time, an investigation of a non-trivial tetrad solution corresponding to Minkowski metric had found one extra dynamical mode [13], though with a somewhat restricted Cauchy data freedom similar to the example of the variable mass Proca field (12). In other words, it rather confirms the results of Ref. [34]
One possible explanation for this strange coincidence is almost obvious [38]. The thing is that the non-trivial tetrad [13], even though being quite non-symmetric, did preserve the condition of ${\mathbb{T}}=0$, while the main mistake of the Hamiltonian analysis [34] was in missing the contributions to Poisson brackets coming from spatial derivatives of the auxiliary scalar field corresponding to $\mathbb{T}$. However, this explanation is then very intriguing in itself. In almost every cosmological setting, away from bounces and the like, the torsion scalar has got a time-like gradient. Then $\mathbb{T}$ can be used precisely as a time variable, and the analysis of Ref. [34] seems to apply perfectly well. If true, it might mean that the model enjoys a preferred foliation, similar to the story of cuscuton fields [39].
@@ 5.3 Conclusions
When exploring a new theoretical idea, we usually start from a Lagrangian density and the variational principle. It gives us very convenient tools for investigation. However, the Hamiltonian formalism is also very important for a variety of reasons. It features very nice symplectic geometry which is good for quantising if we understand anything about quantum physics. And even safely staying inside the classical framework, we happily get a first-order system of time evolution equations. It’s crucial for the whole range of theoretical research, from numerical implementations if needed to rather abstract analyses of mathematical properties.
In the regular situations, the Hamiltonian formalism is relatively well understood, but unfortunately it is very often used as just a recipe from a textbook. However, especially when going for teleparallel geometries, we often encounter ill-posed systems of equations. They might very well be a dead end, but we anyway need to understand what happens with them and how. In my opinion, it is a very interesting and important task, to learn more about it and to see what the Hamiltonian analysis might tell us about such Lagrangian systems.